OpenStudy (anonymous):

Tell me if I am right The length of a rectangle is twice the width. The area is 162^2yds. Find the length and width answer length 162, with 81

OpenStudy (anonymous):

L=2W so L*W=2W*W=162 and\[2W^2=162\]\[W^2=81\]\[W=9\]and the length would be L=2W or L=2(9)=18. check: 18*9=162 as it should

OpenStudy (anonymous):

im not sure I understand

OpenStudy (anonymous):

so the width is 9 and length is 18?

OpenStudy (anonymous):

yes; area=L*W and L=2W so area=(2W)*W or area=2W^2 in general and since area=162 we solve 162=2W^2 after dividing by 2 81=W^2 square root both sides 9=W this is the width; since the length is twice the width (L=2W) we have L=2*9=18 this is the length check: since area=L*W we can check by filling in the given (area=162) and found (W=9; L=18) values 162=9*18 162=162 indicating that these really are the dimensions